OpenStudy (anonymous):

Find the derivative of f(x)=1/4x. I know the answer but I dont understand how to get it.

OpenStudy (anonymous):

1/4

OpenStudy (lastdaywork):

Are you looking for a derivation using limits?

zepdrix (zepdrix):

The function is this:\[\Large\bf\sf \frac{1}{4x}\]yes?

OpenStudy (anonymous):

correct @zepdrix

OpenStudy (anonymous):

-1/x^2

OpenStudy (anonymous):

sry -1/4x^2

OpenStudy (anonymous):

the answer is \[-1/4x^2 \] according to my teacher but I need to know how to get it.

zepdrix (zepdrix):

So what you want to do is use a handy rule of exponents to get the x out of the denominator:\[\Large\bf\sf \frac{1}{4x^1}\quad=\quad \frac{x^{-1}}{4}\quad=\quad \frac{1}{4}x^{-1}\]

zepdrix (zepdrix):

You can apply the power rule from there.

OpenStudy (anonymous):

the rule is

OpenStudy (anonymous):

nx^(n-1)

OpenStudy (anonymous):

your n in this case is negative one cuz when u bring the exponent on the x up to the numerator it becomes negative

OpenStudy (anonymous):

thank you all soo much! i understand now. I just got confused in the beginning steps

OpenStudy (anonymous):

then it would be -1 times (1/4) times x^(-1-1) = (-1)/(4x^2)